# Interpretable Air Pollution Forecasting by Physics-Guided Spatiotemporal Decoupling

### v1

@@ I Introduction
Accurate and reliable urban air pollution forecasts are vital for public health and environmental management. They enable municipal authorities to issue timely health warnings and implement evidence-based traffic management [Crotti2025AURORAE]. The evolution of pollutant concentrations across citywide monitoring networks can be described by the advection-diffusion-reaction equation[seinfeld2016atmospheric]. Two processes dominate: i) Spatial transport, governed mainly by advection along the wind and by near-surface mixing, and ii) Local processes, arising from site-specific emissions and chemical transformation. Existing modeling paradigms struggle to effectively represent and disentangle these processes, leading to a persistent trade-off between predictive accuracy and model interpretability [zhang2025interpretable, zhu2023machine].
Prevailing approaches include numerical and data-driven models. Numerical models provide physical fidelity but are constrained by heavy computation and high-resolution inventories, limiting real-time applications [zhang2024improving]. Deep learning models scale well and achieve strong accuracy [zhang2024mgatt, liang2023airformer, zhang2023meta], yet they compress physical structures into opaque representations with limited diagnostic value. While hybrid physical-neural network designs show promise for improving accuracy [hettige2024airphynet, li2023improving], they often couple training to costly numerical simulators and their attributions remain opaque.
We propose a physics-guided spatiotemporal learning framework that is interpretable-by-design. Our model aligns with physical principles by decoupling the pollutant forecasting into two additive modules with transparent attribution, illustrated in Fig 1 (b) and (c). The first module captures station-wise temporal dependencies using a variant of interpretable attention mechanism. A learnable query attends over historical pollutant and auxiliary-variable sequences, attributing the forecast to specific lags and exogenous variables [zhang2025interpretable]. The second module learns a physics-guided kernel for spatial transport. Its directed weights are dynamically conditioned on meteorological fields and geographic relationships, forming a learnable spatial operator for advection. This separation provides clear physical meaning to each component and enhances both predictive accuracy and interpretability.
@@ II-A Spatiotemporal Forecasting
Numerical air quality models, spanning from regional Chemical Transport Models (CTMs) [cams2017reg] to local street-level models [ottosen2015analysis], provide physical grounding. However, they are computationally intensive and highly sensitive to the accuracy of emission inventories and meteorological inputs, limiting their operational agility in complex urban environments [zhang2023meta]. Data-driven methods, including Recurrent Neural Networks (RNNs), Convolutional Neural Networks (CNNs), Graph Neural Networks (GNNs), and Transformers, excel at modeling complex nonlinear dependencies [wu2020connecting, liang2023airformer, vaswani2017attention]. A persistent challenge, however, is interpretability. Many GNNs employ static, distance-based graphs, which are a poor surrogate for wind-driven atmospheric transport [zhang2024mgatt]. More advanced dynamic graph models often derive connectivity from statistical correlations, resulting in opaque representations that are disconnected from underlying physical processes like advection.
@@ II-B Physics-Guided Models
To build more trustworthy models, researchers are integrating physical knowledge into neural networks. One approach, Physics-Informed Neural Networks (PINNs), uses governing PDEs as soft constraints [raissi2019physics]. This guides the model toward physically plausible solutions but does not render the network architecture inherently interpretable. Another direction involves hybrid models that use neural networks to parameterize components of a physical process, often formulated with Ordinary Differential Equations (ODEs). For instance, AirPhyNet uses a GNN-based differential equation network to model the physical transport of air particles, specifically diffusion and advection[hettige2024airphynet]. While these methods improve physical consistency, their reliance on iterative ODE solvers can introduce significant computational overhead during training. Furthermore, their intricate coupling of neural and physical components can still obscure clear attribution.
@@ III-A Problem Formulation
This input $\mathcal{X}$ contains historical observations and exogenous covariates with varying future availability $m_{i}\leq H$, such as meteorological forecasts and calendar features [zhang2025interpretable].
@@ III-B Governing Equations and Model Formulation
Let $c(\mathbf{x},t)$ be the near-surface PM10 concentration at location $\mathbf{x}$ and time $t$. Its evolution is described by the advection-diffusion-reaction equation [seinfeld2016atmospheric]:
@@
For urban PM10 forecasting, non-local transport between monitoring sites is primarily governed by advection driven by the wind field $\mathbf{u}$ [stull2012introduction]. In contrast, processes such as turbulent mixing ($\mathbf{K}$), chemical reactions ($\mathcal{R}$), and emissions ($S$) operate at a more local scale [zhang2001size, seinfeld2016atmospheric]. This physical scale separation motivates decomposing the PM10 forecast $Y_{s,h}$ at a station $s$ and lead time $h$ into additive advection and local components:
@@ IV-A Embedding of Time-Feature Tokens
We project the $N_{\mathrm{all}}$ time-feature tokens into a $d$-dimensional space and augment them with 1D convolutional and sinusoidal positional embeddings ($\mathbf{P}^{\mathrm{conv1D}}$, $\mathbf{P}^{\mathrm{feat}}$) to form $\mathbf{H}^{(s)}$ [liang2025selection, zhang2023crossformer].
@@
Subsequently, we integrate a learnable station embedding $\mathbf{E}$ via a station-wise affine gating mechanism [chi2023difforecast] to produce the final input representation $\widetilde{\mathbf{H}}\in\mathbb{R}^{S\times N_{\mathrm{all}}\times d}$.
@@ IV-B Station-wise Local Dynamics Encoder
This is achieved through a variant of X2-attention mechanism[zhang2025interpretable]. The keys ($\mathbf{K}$) and values ($\mathbf{V}$) are derived from $\widetilde{\mathbf{H}}$ via linear projections.
@@ V-A Dataset
Our study utilizes hourly air quality data at nine stations from the Air Quality Forecast System in Stockholm, Sweden[engardt2021luften]. The dataset spans from July 2020 to October 2024 and is chronologically partitioned into training (70%), validation (20%), and testing (10%) sets.
To improve predictive accuracy, we incorporate a range of auxiliary variables, including meteorological forecasts, deterministic air quality forecasts from a multi-scale coupled modeling system, and temporal encodings for cyclical patterns. Further details on these auxiliary features and their integration are provided in [zhang2024improving]
@@ V-B Baselines
We evaluate our model against a diverse set of state-of-the-art baselines. These include AirPhyNet [hettige2024airphynet], a hybrid physics-guided model, and Airformer [liang2023airformer], a Transformer architecture tailored for air quality forecasting. We also compare against general-purpose multivariate forecasting models: Crossformer [zhang2023crossformer], which employs two-stage attention, and iTransformer [liu2023itransformer], which inverts the attention mechanism to capture variate-wise correlations. The foundational encoder-only Transformer [vaswani2017attention] serves as a fundamental benchmark. While these models are effective predictors, none are architecturally designed to provide the decoupled spatiotemporal interpretability that our framework offers.
@@ Acknowledgment
The authors thank the Richterska Stiftelsen (2019-00498) and the KTH Centre for Digital Futures (iHorse+ 2024) for supporting this work. The authors also acknowledge Chalmers e-Commons at Chalmers University of Technology, and the National Academic Infrastructure for Supercomputing in Sweden (NAISS) (funded by Swedish Research Council grant 2022-06725) for providing the computational resources.

### v2

@@ I Introduction
Accurate and reliable urban air pollution forecasts are vital for public health and environmental management. They enable municipal authorities to issue timely health warnings and implement evidence-based traffic management [2]. The evolution of pollutant concentrations across citywide monitoring networks can be described by the advection-diffusion-reaction equation[11]. Two processes dominate: i) Spatial transport, governed mainly by advection along the wind and by near-surface mixing, and ii) Local processes, arising from site-specific emissions and chemical transformation. Existing modeling paradigms struggle to effectively represent and disentangle these processes, leading to a persistent trade-off between predictive accuracy and model interpretability [20, 22].
Prevailing approaches include numerical and data-driven models. Numerical models provide physical fidelity but are constrained by heavy computation and high-resolution inventories, limiting real-time applications [19]. Deep learning models scale well and achieve strong accuracy [16, 7, 21], yet they compress physical structures into opaque representations with limited diagnostic value. While hybrid physical-neural network designs show promise for improving accuracy [4, 5], they often couple training to costly numerical simulators and their attributions remain opaque.
We propose a physics-guided spatiotemporal learning framework that is interpretable-by-design. Our model aligns with physical principles by decoupling the pollutant forecasting into two additive modules with transparent attribution, illustrated in Fig 1 (b) and (c). The first module captures station-wise temporal dependencies using a variant of interpretable attention mechanism. A learnable query attends over historical pollutant and auxiliary-variable sequences, attributing the forecast to specific lags and exogenous variables [20]. The second module learns a physics-guided kernel for spatial transport. Its directed weights are dynamically conditioned on meteorological fields and geographic relationships, forming a learnable spatial operator for advection. This separation provides clear physical meaning to each component and enhances both predictive accuracy and interpretability.
@@ II-A Spatiotemporal Forecasting
Numerical air quality models, spanning from regional Chemical Transport Models (CTMs) [12] to local street-level models [9], provide physical grounding. However, they are computationally intensive and highly sensitive to the accuracy of emission inventories and meteorological inputs, limiting their operational agility in complex urban environments [21]. Data-driven methods, including Recurrent Neural Networks (RNNs), Convolutional Neural Networks (CNNs), Graph Neural Networks (GNNs), and Transformers, excel at modeling complex nonlinear dependencies [15, 7, 14]. A persistent challenge, however, is interpretability. Many GNNs employ static, distance-based graphs, which are a poor surrogate for wind-driven atmospheric transport [16]. More advanced dynamic graph models often derive connectivity from statistical correlations, resulting in opaque representations that are disconnected from underlying physical processes like advection.
@@ II-B Physics-Guided Models
To build more trustworthy models, researchers are integrating physical knowledge into neural networks. One approach, Physics-Informed Neural Networks (PINNs), uses governing PDEs as soft constraints [10]. This guides the model toward physically plausible solutions but does not render the network architecture inherently interpretable. Another direction involves hybrid models that use neural networks to parameterize components of a physical process, often formulated with Ordinary Differential Equations (ODEs). For instance, AirPhyNet uses a GNN-based differential equation network to model the physical transport of air particles, specifically diffusion and advection[4]. While these methods improve physical consistency, their reliance on iterative ODE solvers can introduce significant computational overhead during training. Furthermore, their intricate coupling of neural and physical components can still obscure clear attribution.
@@ III-A Problem Formulation
This input $\mathcal{X}$ contains historical observations and exogenous covariates with varying future availability $m_{i}\leq H$, such as meteorological forecasts and calendar features [20].
@@ III-B Governing Equations and Model Formulation
Let $c(\mathbf{x},t)$ be the near-surface PM10 concentration at location $\mathbf{x}$ and time $t$. Its evolution is described by the advection-diffusion-reaction equation [11]:
@@
For urban PM10 forecasting, non-local transport between monitoring sites is primarily governed by advection driven by the wind field $\mathbf{u}$ [13]. In contrast, processes such as turbulent mixing ($\mathbf{K}$), chemical reactions ($\mathcal{R}$), and emissions ($S$) operate at a more local scale [17, 11]. This physical scale separation motivates decomposing the PM10 forecast $Y_{s,h}$ at a station $s$ and lead time $h$ into additive advection and local components:
@@ IV-A Embedding of Time-Feature Tokens
We project the $N_{\mathrm{all}}$ time-feature tokens into a $d$-dimensional space and augment them with 1D convolutional and sinusoidal positional embeddings ($\mathbf{P}^{\mathrm{conv1D}}$, $\mathbf{P}^{\mathrm{feat}}$) to form $\mathbf{H}^{(s)}$ [6, 18].
@@
Subsequently, we integrate a learnable station embedding $\mathbf{E}$ via a station-wise affine gating mechanism [1] to produce the final input representation $\widetilde{\mathbf{H}}\in\mathbb{R}^{S\times N_{\mathrm{all}}\times d}$.
@@ IV-B Station-wise Local Dynamics Encoder
This is achieved through a variant of X2-attention mechanism[20]. The keys ($\mathbf{K}$) and values ($\mathbf{V}$) are derived from $\widetilde{\mathbf{H}}$ via linear projections.
@@ V-A Dataset
Our study utilizes hourly air quality data at nine stations from the Air Quality Forecast System in Stockholm, Sweden[3]. The dataset spans from July 2020 to October 2024 and is chronologically partitioned into training (70%), validation (20%), and testing (10%) sets.
To improve predictive accuracy, we incorporate a range of auxiliary variables, including meteorological forecasts, deterministic air quality forecasts from a multi-scale coupled modeling system, and temporal encodings for cyclical patterns. Further details on these auxiliary features and their integration are provided in [19]
@@ V-B Baselines
We evaluate our model against a diverse set of state-of-the-art baselines. These include AirPhyNet [4], a hybrid physics-guided model, and Airformer [7], a Transformer architecture tailored for air quality forecasting. We also compare against general-purpose multivariate forecasting models: Crossformer [18], which employs two-stage attention, and iTransformer [8], which inverts the attention mechanism to capture variate-wise correlations. The foundational encoder-only Transformer [14] serves as a fundamental benchmark. While these models are effective predictors, none are architecturally designed to provide the decoupled spatiotemporal interpretability that our framework offers.
@@ Acknowledgment
The authors thank the Richterska Stiftelsen (2024-00992) and the KTH Centre for Digital Futures (iHorse+ 2024) for supporting this work. The authors also acknowledge Chalmers e-Commons at Chalmers University of Technology, and the National Academic Infrastructure for Supercomputing in Sweden (NAISS) (funded by Swedish Research Council grant 2022-06725) for providing the computational resources.